OpenStudy (anonymous):

The equation of the common tangent touching the circle (x-3)^2 + y^2 = 9 and the parabola y^2 = 4x above the x axis.

OpenStudy (anonymous):

Let us find the tangent to the parabola \( y = 2 \sqrt x\) at a point \( (m, 2\sqrt m)\) then determine m so this tangent is also tangent to the circle

OpenStudy (anonymous):

It is easy to find it using the slope and a point, you get \[ g(x)=\frac{m+x}{\sqrt{m}} \]

OpenStudy (anonymous):

Now substitute the point (x,g(x)) in the circle to find the point of intersections of the tangent and the circle, you get \[ \left(\frac{m+x}{\sqrt{m}}\right)^2+(x-3)^2-9=\\\frac{x ^2}{m}+m+x^2-4 x=\\\left(\frac{1}{m}+1\right) x^2-4 x + m=0 \]

OpenStudy (anonymous):

Now \[ \left(\frac{1}{m}+1\right) x^2-4 x + m=0 \] is a quadratic in x. To be tangent, we need the two roots to be equal, so we need \[b^2-4 a c=0=16-4 \left(\frac{1}{m}+1\right) m=12-4 m=0 \] Hence m=3

OpenStudy (anonymous):

Finally the equation of the tangent is \[ g(x)=\frac{m+x}{\sqrt{m}}=\frac{3+x}{\sqrt{3}} \]

OpenStudy (anonymous):

In the attached graph the red line is the common tangent and the blue graph is that of the parabola. @satellite73 @experimentX @ganeshie8

OpenStudy (anonymous):

May be the people, I tagged above can do it in an easier way.

OpenStudy (anonymous):

 Your answer is correct. But what's wrong with my method?When I equate and square the second terms of both. Why doesn't that give the correct ans?